# Next-Embedding Prediction Makes Strong Vision Learners

### v1

@@ 2 Related Work
Generative pretraining has driven major advances in language [radford2018gpt, radford2019gpt2, brown2020gpt3, achiam2023gpt4] and has a long visual lineage. Early pixel-level generative pretraining (iGPT) showed transferable features but struggled with very long sequences and weak semantic alignment [chen2020igpt]. Subsequent work introduced discrete tokenizers [van2017vqvae, esser2021vqgan, lee2022rqvae, sun2024llamagen, pang2024randar] to compress images into short code sequences amenable to transformer generative pretraining [esser2021vqgan, ramesh2021dalle, lee2022rqvae], enabling large-scale image generation. Recent scaling efforts, such as VAR [tian2024var] and LlamaGen [sun2024llamagen], further close or surpass the diffusion modeling performance. In parallel, tokenizer-free generative pretraining operates directly in continuous spaces [fan2024fluid, li2024mar], and masked autoregression blends bidirectional context with ordered generation for efficiency and quality [li2024mar, wu2025dcar, chang2022maskgit, li2023mage]. Unlike these works, we treat prediction itself as the learning signal and operate directly in the embedding space without a generative decoder.
@@ 6 Conclusion and Future Work
Recent Large Language Models [touvron2023llama] increasingly adopt tied embeddings, where the input and output embedding matrices are shared. This mechanism is effectively equivalent to predicting the next embedding in the latent space, which is precisely the principle underlying our framework. From this perspective, our work does not introduce a new paradigm, but rather reveals a unifying view: different modalities can be trained under the same objective, with embeddings serving as a common representational currency.
@@ Acknowledgment
This work is supported in part by a departmental fellowship to Sihan Xu and Weinberg Cognitive Science Fellowship to Sihan Xu and Ziqiao Ma. We acknowledge the computing resources of Google Cloud Platform and Lambda Labs. The authors would like to thank Jiayi Pan and Freda Shi for their valuable feedback.
@@ Appendix B Methodology Comparisons
Relation to GPT [radford2018gpt, chen2020igpt].
NEPA shares the causal prediction idea of GPT, but works fully in the continuous embedding space. We do not use task-specific decoders, tokenizers, or language heads during pretraining.
@@ Appendix C Implementation Details
We first prototyped and validated NEPA using the timm [rw2019timm] library. For large-scale experiments, we switched to the Hugging Face ecosystem, relying on transformers [wolf-etal-2020-transformers], datasets [lhoest-etal-2021-datasets], and evaluate. Our implementation is mainly based on the ViT code in transformers.models.vit.modeling_vit, and the training scripts are adapted from the official run_mim and run_image_classification examples. For data augmentation, we use the timm utilities, in particular Mixup and create_transform. All ImageNet experiments use the ILSVRC/imagenet-1k dataset accessed through Hugging Face datasets. ADE20K experiments follow the standard recipes and data processing pipelines from mmsegmentation [mmseg2020]. Our code and model is available at https://sihanxu.github.io/nepa.

### v2

@@ 2 Related Work
Generative pretraining has driven major advances in language [radford2018gpt, radford2019gpt2, brown2020gpt3, achiam2023gpt4] and has a long visual lineage. iGPT autoregresses over discrete pixel tokens and learns transferable features, but suffers from long sequences and weak semantic alignment [chen2020igpt]. AIM series [el2024scalable, fini2025multimodal] extend to the continuous counterpart, supervised on predefined pixel-space targets as continuous patch values. Subsequent work introduced discrete tokenizers [van2017vqvae, esser2021vqgan, lee2022rqvae, sun2024llamagen, pang2024randar] to compress images into short code sequences amenable to transformer generative pretraining [esser2021vqgan, ramesh2021dalle, lee2022rqvae], enabling large-scale image generation. Recent scaling efforts, such as VAR [tian2024var] and LlamaGen [sun2024llamagen], further close or surpass the diffusion modeling performance. In parallel, tokenizer-free generative pretraining operates directly in continuous spaces [fan2024fluid, li2024mar], and masked autoregression blends bidirectional context with ordered generation for efficiency and quality [li2024mar, wu2025dcar, chang2022maskgit, li2023mage]. Unlike these works, we treat prediction itself as the learning signal and operate directly in the embedding space without a generative decoder.
@@ 6 Conclusion and Future Work
Recent Large Language Models increasingly adopt tied embeddings, where the input and output embedding matrices are shared. This mechanism is effectively equivalent to predicting the next embedding in the latent space, which is precisely the principle underlying our framework. From this perspective, our work does not introduce a new paradigm, but rather reveals a unifying view: different modalities can be trained under the same objective, with embeddings serving as a common representational currency.
@@ Acknowledgment
This work is supported in part by a departmental fellowship to Sihan Xu and Weinberg Cognitive Science Fellowship to Sihan Xu and Ziqiao Ma. We acknowledge the computing resources of Google Cloud Platform and Lambda Labs. We sincerely thank Steven Bucaille for opening a pull request that adds NEPA to the transformers ecosystem. We thank Jiayi Pan and Freda Shi for their valuable feedback.
@@ Appendix B Methodology Comparisons
Relation to GPT [radford2018gpt], iGPT [chen2020igpt] and AIM [el2024scalable, fini2025multimodal].
NEPA shares the causal prediction idea, but works fully in the continuous embedding space. We do not use task-specific decoders, tokenizers, or language heads during pretraining. iGPT and AIM-style methods supervise against predefined pixel-space targets (discrete pixel tokens or continuous patch values), whereas NEPA supervises in representation space via embedding similarity, a learned, moving target defined by the embedding rather than pixels.
@@ Appendix C Implementation Details
We first prototyped and validated NEPA using the timm [rw2019timm] library. For large-scale experiments, we switched to the Hugging Face ecosystem, relying on transformers [wolf-etal-2020-transformers], datasets [lhoest-etal-2021-datasets], and evaluate. Our implementation is mainly based on the ViT code in transformers.models.vit.modeling_vit, and the training scripts are adapted from the official run_mim and run_image_classification examples. For data augmentation, we use the timm utilities, in particular Mixup and create_transform. All ImageNet experiments use the ILSVRC/imagenet-1k dataset accessed through Hugging Face datasets. ADE20K experiments follow the standard recipes and data processing pipelines from mmsegmentation [mmseg2020].Our code and model is available at https://github.com/SihanXU/nepa.